# Strong-coupling quantum thermodynamics using a superconducting flux qubit

### v2

@@ I Introduction
In this work, we report on two devices: Device 1 and Device 2. Device 1 is a thermal device architecture that forms the core of the reported experiment. Device 2, which plays a supporting role, is a spectroscopy-enabled device designed for performing qubit spectroscopy useful for complete understanding of the experiment. In Device 1, we experimentally investigate quantum heat transport between two reservoirs mediated by an interconnected artificial atom, which is a superconducting flux qubit. The concept of Device 1, illustrated in Fig. 1 (a), depicts photons energized by Joule heating in a metallic reservoir transferring heat to a cooler region, but being blocked by a superconducting qubit, which, modulated by a weak magnetic field, acts as a controlled gate, allowing photon passage only at certain flux values. By manipulating the quantum states of the qubit via a tunable, small external magnetic field, we demonstrate controllable heat transport across it [6, 47]. The observed heat transport exhibits a triplet structure, with the central peak reaching its maximum at the degeneracy point, corresponding to half a flux quantum in a superconducting flux qubit, as depicted in Fig. 1 (c). In an RF environment, utilizing a read-out resonator, characteristic transmission across two ports of a spectroscopy enabled device shown in panel (b) of Fig. 1, is observed upon changing the magnetic field, as observed in Fig. 1 (c). By overlaying these two signals, a preliminary understanding of the system’s behavior under varying magnetic field in both DC and RF environments is achieved, revealing the origin of the satellite peaks in the triplet heat transport signal when the qubit frequency aligns with the resonator frequency, outlined in Fig. 1 (c). In Ref. [6], a weakly-anharmonic transmon qubit was used to demonstrate the concept of such a measurement, however the device performance was limited by its weak qubit-resonator ($g$) and resonator-reservoir ($\gamma$) coupling strengths. The reported flux qubit architecture enables access to ultra-strong and even deep-ultra-strong coupling regimes with minimal modifications to the qubit design. These regimes, which remain experimentally unexplored in the field of quantum thermodynamics, are largely inaccessible using transmon-type qubit architectures. The novelty of our approach lies in exploiting the length-dependent mutual inductance of the coupling element, which enables further enhancement of the qubit–resonator coupling strength, pushing it into the ultra-strong coupling regime and beyond [29, 26, 28, 30]. A theoretical perspective towards high performance of such devices is addressed in Ref. [18], where the importance of couplings $g$ and $\gamma$ is discussed. In this work, within the framework of strong coupling between the reservoir, resonator and qubit, our results demonstrate an order of magnitude increase in power modulation along with a nearly 100% on-off switching ratio of heat current in the photon dominated heat transport regime. With a primary focus on experimental observations, this work also introduces a theoretical model that quantitatively captures the observed heat current and qualitatively explains the flux-dependent power. The device design is reported in Fig. 2, with detailed description provided in Section II of this article.
@@
Figure 2: Device design and circuit. (a) An optical image of the reported device. At the center, within the solid black box, is a superconducting qubit coupled to Resonator 1 (green) and Resonator 2 (blue) and connected to the ground plane. The lower sections display Reservoir 1 and Reservoir 2, along with a brief measurement scheme provided in detail in the article Appendix B. (b) Conceptual schematic showing both reservoirs, each connected to a phonon bath at temperature $T_{0}$, with thermal conductance $G_{ep,1}$ and $G_{ep,2}$. The qubit’s flux-dependent frequency $(f(\Phi))$ controls heat transport ($Q_{G\Phi}$), acting as a tunable quantum heat switch. (c) Circuit diagram shows a resistor $(R)$ connected to a half-wavelength resonator and qubit, with $i_{n}$ as current noise and $\Phi=Mi(t)$ for input coil inductance. Reported SEM images have been colored for clarity.
@@ II Device design
Since the primary experiment is performed using Device 1, we focus on its design in this section. The design details of Device 2 are discussed separately in Section IV. The hybrid quantum system is composed of a flux qubit galvanically connected to two nominally identical superconducting half-wavelength coplanar resonators with designed frequency of $f_{r}\approx 6.5$ GHz, which are in turn coupled to two copper reservoirs referred to as Source (S) and Drain (D) reservoir. Thus, the device architecture can be succinctly described as: Source – Resonator 1 – Flux qubit – Resonator 2 – Drain. For convenience, we will refer to the Source reservoir as Reservoir 1 and the Drain reservoir as Reservoir 2 throughout the article. The measured resistance of the reservoirs $R$ are $\approx 6~{}\Omega$ at helium temperatures. The measured device is shown in Fig.  2 (a). The flux qubit in the middle is a superconducting loop interrupted by three Josephson junctions in series where two of them are designed to be identical with the critical current $I_{C2}=I_{C3}=I_{C}$, while the area of the first junction is reduced by a factor of $\alpha\approx 0.6$ with critical current $I_{C1}=\alpha I_{C}$. The scanning electron microscope (SEM) image of the zoom-in of the flux qubit in Fig.  2 (a) shows the galvanic connections to Resonator 1 and Resonator 2 via the inductances $L_{1}$ and $L_{2}$ realized by a thin superconducting Al wire, facilitating couplings $g_{1}$ and $g_{2}$ with each resonator, respectively. The coupling constants are proportional to the corresponding inductances as $g_{j}\propto L_{j}$ (where $L_{1}\approx L_{2}$[26, 27]. The fundamental frequencies of the resonators and the qubit-resonator coupling have been carefully designed to be identical on both sides of the qubit (side 1 and side 2) to maintain symmetry in the device. The estimated coupling strength $g_{1}\simeq g_{2}\simeq 200$ MHz, was extracted by fitting spectroscopy data described in section IV.
@@
Figure 3: Flux-dependent power transport. (a) Flux-dependent power transport, $P_{1\rightarrow 2}$, across the qubit, driven by a magnetic flux generated by the coil, with power from Reservoir 1 (red) absorbed by Reservoir 2 (blue), at $T_{0}=80\,$mK. The bottom plot shows the modulation of Reservoir 2’s temperature (left y-axis) and transported power (right y-axis) as a function of the applied magnetic field. The region within the dashed rectangle in the center of the image is enlarged in panels (b) and (c). In panel (b), the y-axis represents temperature, while in panel (c), it represents the power absorbed by Reservoir 2. (b) $T_{2}$ vs applied magnetic flux at different $T_{1}$. (c) $P_{1\to 2}$ vs applied magnetic flux at different $T_{1}$. (d) Averaged data (pink curve) from multiple sweeps shows small satellite peaks (black arrows). (e) $P_{1\to 2}$ from the theoretical model. (f) Satellite peaks captured by the theoretical model.
Figure 4: Device spectroscopy. (a) Two-tone spectroscopy measurement probing the interaction of a strongly coupled qubit–resonator system via a read-out resonator at the readout frequency $f_{\text{read-out}}\approx 7\,\text{GHz}$. The estimated coupling strength is $2g\approx 390\,\text{MHz}$, obtained by fitting with the J-C model. (b) Qubit energy spectrum vs applied magnetic flux, with minimum frequency ($\approx 4.0$ GHz) at half-flux. The orange dashed-line shows the J-C model fit; the white dashed-line fits $\hbar\,\epsilon=2I_{p}\,(\Phi-\Phi_{0}/2)$, with $I_{p}\approx 21$ nA. The avoided crossing with a stationary feature around $\approx 4.4~{}\mathrm{GHz}$ arises from the interaction of the system with a parasitic resonance which is disregarded in the analysis of the results. (c) Dominant transitions from the theoretical model responsible for heat transport across the qubit.
@@ IV Qubit read-out
Here, we report on Device 2 that enables us to perform rf spectroscopy on it to extract parameters such as the resonator frequency, qubit transition frequency and their coupling strengths. Since the resonator frequency and coupling strength are merely geometry dependent quantities, the spectroscopy replica gives estimates for these parameters. Yet the qubit frequency is more susceptible to precise fabrication processes, and therefore variations between Device 1 and 2 may be large. These parameters are essential for supporting the interpretation of the thermal transport measurements performed using Device 1. The Device 2 design schematic is shown in Fig. 5. It features a signal feed-line that is capacitively coupled to one end of a quarter-wavelength read-out resonator, with a designed read-out frequency $f_{read-out}=$ 7.07.07.0~{}7.0GHz, see panel (b) of Fig. 5. The opposite end of the read-out resonator is positioned in close proximity to the flux qubit and is inductively terminated to the on-chip ground plane, as shown in Fig. 5 (c). As reported in Ref. [49], the broadening of the resonance line-width is expected with copper terminations. Therefore, for measurement convenience, in the spectroscopy device, copper terminations are not used. Instead, a superconducting aluminum bridge is employed to connect the resonator to the niobium ground plane, see panels (d) and (e) of Fig. 5. We use continuous wave transmission coefficient measurements to probe the device characteristics.
@@
Details about the cryostat and the measurement setup are covered in Appendix B of this article and they remain similar to the scheme employed in Ref. [29]. To characterize the qubit properties we perform single and two-tone spectroscopy. The weakly coupled qubit-read-out resonator interaction is detected by the change in the read-out resonator frequency as a function of applied magnetic field. The qubit energy spectra as a function of applied magnetic flux is measured using two-tone spectroscopy where the second-tone is used to excite the qubit and the probe tone detects the change in qubit state. For more details on single and two-tone spectroscopy techniques see Ref. [29]. In this article, the two-tone qubit spectroscopy data is presented in panels (a) and (b) of Fig. 4. The splitting visible in panel (a) highlights the strong interaction $(2g)$ between the qubit frequency and the resonances observed in the device, while panel (b) reveals the qubit energy spectrum, which eventually reaches a minimum at approximately 3.9 GHz at qubits’s degeneracy point. To estimate the qubit-resonator interaction energy $(2g)$ and qubit’s persistent current $(I_{p})$ of the device we fit the plots shown in panel (a) and (b) of Fig. 4, with well-known Jaynes-Cummings model (J-C) [50] and $\hbar\,\epsilon=2\,I_{p}\,(\,\Phi\,-\,\Phi_{0}/2\,)$ [51], where $\omega=\sqrt{\epsilon^{2}+\Delta^{2}}$. In particular, we may write $hf_{q}(\Phi)=hf_{q}^{0}\sqrt{1+\left(\frac{I_{p}}{hf_{q}^{0}}\right)^{2}\left(\frac{\Phi}{\Phi_{0}}-\frac{1}{2}\right)^{2}}$, where $f_{q}(\Phi)$ and $f_{q}^{0}$ represent the qubit frequencies at flux $\Phi$ and at $\Phi=0$, respectively. For this particular sample, we obtain $I_{p}=21~{}\mathrm{nA}$.
@@
[…]
@@
where $|\psi_{k}\rangle=\sum_{i=0}^{3}a_{ki}|i\rangle$, with the coefficients $a_{ki}$ provided in Appendix 3. Here, $S_{\Phi}(\omega_{ij})$ represents the spectral density of flux noise, which will be discussed later in this section. Further details on ^$\partial\hat{\mathcal{H}}/{\partial\Phi}$ and the matrix elements $|\langle\psi_{i}|{\partial\hat{\mathcal{H}}}/{\partial\Phi}|\psi_{j}\rangle|^{2}$ can be found in Appendix 3 of the article.
@@
[…]
@@
Figures 3 (a-d) and 3 (f) show the flux dependence of the power in the experiment and that obtained by the model above, respectively. The parameters used in the calculations and plotting the theoretical curves are as follows: $R=6\,\Omega$, $L=0.8\,$nH, $I_{p}=30\,$nA, the minimum energy of the qubit at half flux, $f_{\rm q}^{(0)}=2\,$GHz , and $f_{\rm r}=7\,$GHz. Our model has some simplifying assumptions. We restrict it to a single-excitation basis and use the Fermi golden rule transition rates. These formulations are fine as long as temperatures are sufficiently low and couplings are not too strong.

### v3

@@ I Introduction
In this work, we report on two devices: Device 1 and Device 2. Device 1 is a thermal device architecture that forms the core of the reported experiment. Device 2, which plays a supporting role, is a spectroscopy-enabled device designed for performing qubit spectroscopy useful for complete understanding of the experiment. In Device 1, we experimentally investigate quantum heat transport between two reservoirs mediated by an interconnected artificial atom, which is a superconducting flux qubit. The concept of Device 1, illustrated in Fig. 1 (a), depicts photons energized by Joule heating in a metallic reservoir transferring heat to a cooler region, but being blocked by a superconducting qubit, which, modulated by a weak magnetic field, acts as a controlled gate, allowing photon passage only at certain flux values. By varying the external magnetic field, we tune the qubit frequency accordingly with respect to that of the cavity, which allows us to control heat transport across the qubit [6, 47]. The observed heat transport exhibits a triplet structure, with the central peak reaching its maximum at the degeneracy point, corresponding to half a flux quantum in a superconducting flux qubit, as depicted in Fig. 1 (c). In an RF environment, utilizing a read-out resonator, characteristic transmission across two ports of a spectroscopy enabled device shown in panel (b) of Fig. 1, is observed upon changing the magnetic field, as observed in Fig. 1 (c). By overlaying these two signals, a preliminary understanding of the system’s behavior under varying magnetic field in both DC and RF environments is achieved, revealing the origin of the satellite peaks in the triplet heat transport signal when the qubit frequency aligns with the resonator frequency, outlined in Fig. 1 (c). In Ref. [6], a weakly-anharmonic transmon qubit was used to demonstrate the concept of such a measurement, however the device performance was limited by its weak qubit-resonator ($g$) and resonator-reservoir ($\gamma$) coupling strengths. The reported flux qubit architecture enables access to ultra-strong and even deep-ultra-strong coupling regimes with minimal modifications to the qubit design. These regimes, which remain experimentally unexplored in the field of quantum thermodynamics, are largely inaccessible using transmon-type qubit architectures. The novelty of our approach lies in exploiting the length-dependent mutual inductance of the coupling element, which enables further enhancement of the qubit–resonator coupling strength, pushing it into the ultra-strong coupling regime and beyond [29, 26, 28, 30]. A theoretical perspective towards high performance of such devices is addressed in Ref. [18], where the importance of couplings $g$ and $\gamma$ is discussed. In this work, within the framework of strong coupling between the reservoir, resonator and qubit, our results demonstrate an order of magnitude increase in power modulation along with a nearly 100% on-off switching ratio of heat current in the photon dominated heat transport regime. With a primary focus on experimental observations, this work also introduces a theoretical model that quantitatively captures the observed heat current and qualitatively explains the flux-dependent power. The device design is reported in Fig. 2, with detailed description provided in Section II of this article.
@@
Figure 2: Device design and circuit. (a) An optical image of the reported device. At the center, within the solid black box, is a superconducting qubit coupled to Resonator 1 (green) and Resonator 2 (blue) and connected to the ground plane. The lower sections display Reservoir 1 and Reservoir 2, along with a brief measurement scheme provided in detail in the article Appendix B. (b) Conceptual schematic showing both reservoirs, each connected to a phonon bath at temperature $T_{0}$, with thermal conductance $G_{ep,1}$ and $G_{ep,2}$, respectively. The qubit’s flux-dependent frequency $(f(\Phi))$ controls heat transport ($Q_{G\Phi}$), thus it acts as a tunable quantum heat switch. (c) Circuit diagram shows a resistor $(R)$ connected to a half-wavelength resonator and qubit, with $i_{n}$ as current noise and $\Phi=Mi(t)$ for input coil inductance. Reported SEM images have been colored for clarity.
@@ II Device design
Since the primary experiment is performed using Device 1, we focus on its design in this section. The design details of Device 2 are discussed separately in Section IV. The hybrid quantum system is composed of a flux qubit galvanically connected to two nominally identical superconducting half-wavelength coplanar resonators with designed frequency of $f_{r}\approx 6.5$ GHz, which are in turn coupled to two copper reservoirs referred to as Source (S) and Drain (D) reservoir. Thus, the device architecture can be succinctly described as: Source – Resonator 1 – Flux qubit – Resonator 2 – Drain. For convenience, we will refer to the Source reservoir as Reservoir 1 and the Drain reservoir as Reservoir 2 throughout the article. The measured resistance of the reservoirs $R$ are $\approx 6~\Omega$ at helium temperatures. The reservoirs are coupled to their respective resonators with a coupling strength given by $\gamma/2\pi=f_{r}/Q_{R}$ $\approx 500$ MHz. Here, $Q_{R}$ is the quality factor of the reservoir given by $Q_{R}=\frac{\pi}{2}Z_{0}/R$, where $Z_{0}(\approx 50\Omega)$ is the designed characteristic impedance of the resonator [6]. The measured device is shown in Fig.  2 (a). The flux qubit in the middle is a superconducting loop interrupted by three Josephson junctions in series where two of them are designed to be identical with the critical current $I_{C2}=I_{C3}=I_{C}$, while the area of the first junction is reduced by a factor of $\alpha\approx 0.6$ with critical current $I_{C1}=\alpha I_{C}$. The scanning electron microscope (SEM) image of the zoom-in of the flux qubit in Fig.  2 (a) shows the galvanic connections to Resonator 1 and Resonator 2 via the inductances $L_{1}$ and $L_{2}$ realized by a thin superconducting Al wire, facilitating couplings $g_{1}$ and $g_{2}$ with each resonator, respectively. The coupling constants are proportional to the corresponding inductances as $g_{j}\propto L_{j}$ (where $L_{1}\approx L_{2}$[26, 27]. The fundamental frequencies of the resonators and the qubit-resonator coupling have been carefully designed to be identical on both sides of the qubit (side 1 and side 2) to maintain symmetry in the device. The estimated coupling strength $g_{1}/2\pi\simeq g_{2}/2\pi\simeq 200$ MHz, was extracted by fitting spectroscopy data described in section IV.
@@
Figure 3: Flux-dependent power transport. (a) Flux-dependent power transport, $P_{1\rightarrow 2}$, across the qubit, driven by a magnetic flux generated by the coil, with power from Reservoir 1 (red) absorbed by Reservoir 2 (blue), at $T_{0}=80\,$mK. The bottom plot shows the modulation of Reservoir 2’s temperature (left y-axis) and transported power (right y-axis) as a function of the applied magnetic field. The region within the dashed rectangle in the center of the image is enlarged in panels (b) and (c). In panel (b), the y-axis represents temperature, while in panel (c), it represents the power absorbed by Reservoir 2. (b) $T_{2}$ vs applied magnetic flux at different $T_{1}$. (c) $P_{1\to 2}$ vs applied magnetic flux at different $T_{1}$. (d) Averaged data (pink curve) from multiple sweeps shows small satellite peaks (black arrows). (e) $P_{1\to 2}$ from the theoretical model. (f) Satellite peaks captured by the theoretical model. The average RMS noise, calculated over the range $\pm 0.25\,\Phi/\Phi_{0}$ for the data shown in panel (a), is $\approx$~0.14 fW, which is considerably smaller than the amplitude of the signal of interest.
Figure 4: Device spectroscopy. (a) Two-tone spectroscopy measurement probing the interaction of a strongly coupled qubit–resonator system via a read-out resonator at the readout frequency $f_{\text{read-out}}\approx 7\,\text{GHz}$. The estimated coupling strength is $g/2\pi\approx 200\,\text{MHz}$, obtained by fitting with the J-C model. (b) Qubit energy spectrum vs applied magnetic flux, with minimum frequency ($\approx 4.0$ GHz) at half-flux. The orange dashed-line shows the J-C model fit; the white dashed-line fits $\hbar\,\epsilon=2I_{p}\,(\Phi-\Phi_{0}/2)$, with $I_{p}\approx 21$ nA. The avoided crossing with a stationary feature around $\approx 4.4~\mathrm{GHz}$ arises from the interaction of the system with a parasitic resonance which is disregarded in the analysis of the results. (c) Dominant transitions from the theoretical model responsible for heat transport across the qubit.
@@ IV Qubit read-out
Here, we report on Device 2 that enables us to perform rf spectroscopy on it to extract parameters such as the resonator frequency, qubit transition frequency and their coupling strengths. Since the resonator frequency and coupling strength are merely geometry dependent quantities, the spectroscopy replica gives estimates for these parameters. Yet the qubit frequency is more susceptible to precise fabrication processes, and therefore variations between Device 1 and 2 may be large. These parameters are essential for supporting the interpretation of the thermal transport measurements performed using Device 1. The Device 2 design schematic is shown in Fig. 5. It features a signal feed-line that is capacitively coupled to one end of a quarter-wavelength read-out resonator, with a designed read-out frequency $f_{read-out}=$ 7.07.0~GHz, see panel (b) of Fig. 5. The opposite end of the read-out resonator is positioned in close proximity to the flux qubit and is inductively terminated to the on-chip ground plane, as shown in Fig. 5 (c). As reported in Ref. [49], the broadening of the resonance line-width is expected with copper terminations. Therefore, for measurement convenience, in the spectroscopy device, copper terminations are not used. Instead, a superconducting aluminum bridge is employed to connect the resonator to the niobium ground plane, see panels (d) and (e) of Fig. 5. We use continuous wave transmission coefficient measurements to probe the device characteristics.
@@
Details about the cryostat and the measurement setup are covered in Appendix B of this article and they remain similar to the scheme employed in Ref. [29]. To characterize the qubit properties we perform single and two-tone spectroscopy. The weakly coupled qubit-read-out resonator interaction is detected by the change in the read-out resonator frequency as a function of applied magnetic field. The qubit energy spectra as a function of applied magnetic flux is measured using two-tone spectroscopy where the second-tone is used to excite the qubit and the probe tone detects the change in qubit state. For more details on single and two-tone spectroscopy techniques see Ref. [29]. In this article, the two-tone qubit spectroscopy data is presented in panels (a) and (b) of Fig. 4. The splitting visible in panel (a) highlights the strong interaction $(2g)$ between the qubit frequency and the resonances observed in the device, while panel (b) reveals the qubit energy spectrum, which eventually reaches a minimum at approximately 3.9 GHz at qubits’s degeneracy point. To estimate the qubit-resonator interaction energy and qubit’s persistent current $(I_{p})$ of the device we fit the plots shown in panel (a) and (b) of Fig. 4, with well-known Jaynes-Cummings model (J-C) [50] and $\hbar\,\epsilon=2\,I_{p}\,(\,\Phi\,-\,\Phi_{0}/2\,)$ [51], where $\omega=\sqrt{\epsilon^{2}+\Delta^{2}}$. In particular, we may write $hf_{q}(\Phi)=hf_{q}^{0}\sqrt{1+\left(\frac{I_{p}}{hf_{q}^{0}}\right)^{2}\left(\frac{\Phi}{\Phi_{0}}-\frac{1}{2}\right)^{2}}$, where $f_{q}(\Phi)$ and $f_{q}^{0}$ represent the qubit frequencies at flux $\Phi$ and at $\Phi=0$, respectively. For this particular sample, we obtain $I_{p}=21~\mathrm{nA}$.
@@
$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\rm q}+\sum_{i=1,\,2}\hat{\mathcal{H}}_{\rm i}+\sum_{i=1,\,2}\hat{\mathcal{H}}_{\rm qci}+\hat{\mathcal{H}}_{\rm 12}.$ (1)
Here $\hat{\mathcal{H}}_{\rm q}=hf_{\rm q}(\Phi)\hat{b}^{\dagger}\hat{b}$ is the Hamiltonian of the qubit, $\hat{\mathcal{H}}_{\rm i}=\sum_{i=1,\,2}hf_{i}\hat{a}_{i}^{\dagger}\hat{a}_{i}$ that of each resonator, and we approximate the coupling between the resonator $i$ and the qubit as $\hat{\mathcal{H}}_{\rm qci}=-\sum_{i=1,\,2}hg_{i}(\hat{a}_{i}+\hat{a}_{i}^{\dagger})(\hat{b}+\hat{b}^{\dagger})$, and that between the two resonators as $\hat{\mathcal{H}}_{\rm 12}=-h\gamma(\hat{a}_{1}+\hat{a}_{1}^{\dagger})(\hat{a}_{2}+\hat{a}_{2}^{\dagger})$ [48]. Here, $\hat{b}^{\dagger}$ and $\hat{b}$ represent the creation and annihilation operators for the qubit, and similarly, $\hat{a}^{\dagger}$ and $\hat{a}$ are for resonator $i$, $g_{i}$ denotes the coupling constant between the qubit and resonator $i$, and $\gamma$ between the two resonators. We use the Hilbert space $\big\{|0\rangle\equiv|g00\rangle$, $|1\rangle\equiv|e00\rangle$, $|2\rangle\equiv|g10\rangle$ and $|3\rangle\equiv|g01\rangle\big\}$, where the three entries in each state refer to the qubit which is either in the ground state $|g\rangle$ or excited state $|e\rangle$, Resonator 1, and Resonator 2, respectively. The non-vanishing transitions between the eigenstates, which contribute to the heat transport in our device, are spectroscopically examined and presented in panels (a) and (b) of Fig. 4. These transitions are modeled as illustrated in panel (c) of the same figure.
The transition rate induced by the reservoir r(r=1, 2)r~(r=1,\,2) between the normalized eigenstates $|\psi_{i}\rangle$ and $|\psi_{j}\rangle$ of Hamiltonian (7) reads
@@
where $|\psi_{k}\rangle=\sum_{i=0}^{3}a_{ki}|i\rangle$, with the coefficients $a_{ki}$ provided in Appendix 3. Here, $S_{\Phi}(\omega_{ij})$ represents the spectral density of flux noise, which will be discussed later in this section. Further details on $\partial\hat{\mathcal{H}}/{\partial\Phi}$ and the matrix elements $|\langle\psi_{i}|{\partial\hat{\mathcal{H}}}/{\partial\Phi}|\psi_{j}\rangle|^{2}$ can be found in Appendix 3 of the article.
@@
Here $\Gamma_{i\rightarrow j}=\Gamma_{i\rightarrow j}^{r=1}+\Gamma_{i\rightarrow j}^{r=2}$. The power to reservoir $r$, (r=1, 2r=1,\,2) then reads
@@
Figures 3 (a-d) and 3 (f) show the flux dependence of the power in the experiment and that obtained by the model above, respectively. The parameters used in the calculations and plotting the theoretical curves are as follows: $R=6\,\Omega$, L=0.8L=0.8\,nH, $I_{p}=30\,$nA, the minimum energy of the qubit at half flux, $f_{\rm q}^{(0)}=2\,$GHz , and $f_{\rm r}=7\,$GHz. Our model has some simplifying assumptions. We restrict it to a single-excitation basis and use the Fermi golden rule transition rates. These formulations are fine as long as temperatures are sufficiently low and couplings are not too strong.